# Beyond Patches: Mining Interpretable Part-Prototypes for Explainable AI

### v1

@@ 3.3. Step 3: Concept Activation Mining
(5) $\mathbf{z}^{i}=[\mathcal{S}(C(l,p,j),\mathbf{f}_{p}^{i})\;\forall p\in\{1..K\}\;\text{and}\;j\in\{1..L\}]\in\mathbb{R}^{d_{c}}$
where $\mathcal{S}$ is cosine distance and [.][.][ . ] represents the vector concatenation. $d_{c}$ is the number of total concept centroids. We aim to explain each output class of our model using a small set of interpretable concepts. We, therefore, train a sparse classifier on top of $z^{i}$ to obtain the final classification scores $o^{i}=W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b$ and the predicted class $\hat{y}^{i}=\arg\max(o^{i})$. Here, $W_{1}\in\mathbb{R}^{d_{c}\times L}$, $W_{2}\in\mathbb{R}^{d_{f}\times L}$ and $b\in\mathbb{R}^{L}$ denote the classification weights and bias term, respectively. This layer is trained with the following classification loss, using the GLM-SAGA optimizer (Wong et al., 2021):
(6) $\mathcal{L}_{\text{cl}}=\mathcal{L}_{ce}(W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b,y^{i})+\lambda\mathcal{R}(W_{1}),$
where $\mathcal{L}_{\text{ce}}$ is the cross-entropy loss, $y^{i}$ is the label, $\lambda$ is the sparsity regularization strength and $\mathcal{R}(W)=(1-\gamma)\frac{1}{2}\|W\|_{F}+\gamma\|W\|_{1}$, here $\|W\|_{F}$ is the Forbenius norm and $\|W\|_{1}$ is the element-wise matrix.
Figure 4. Concept Mining Module. First, a part centroid list is generated by applying DBSCAN(Ester et al., 1996) clustering within each class and computing the centers of the resulting clusters. These centroids serve as frequently occurring concept prototypes. Next, to activate the most relevant concept for a given part feature, the similarity (inverse distance) to the centroids is measured. These values are then used as CAVs to inform the model’s decision-making process .
3.4. Overall Training
We jointly optimize the network in an end-to-end manner by loss:
(7) $\mathcal{L}=\mathcal{L}_{\text{part}}+\beta\mathcal{L}_{\text{cl}},$
where $\beta$ is hyperparameter. For the first step of training, we set $\beta=0$; then, once the part-prototypical concepts are learned, we set $\beta=2$ based on our hyperparameter analysis.
4. Experiments and Results
@@
Stanford Cars Resnet50+Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 0.16 - 0.23 22.74 65.4 84.1
Resnet50+W $\times$ Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 1.54 - 6.61 22.74 65.4 84.1
ProtoPNet + Act 45.70$\pm$4.9 11.22$\pm$5.5 2.2-55.64 60.92 69.6 84.5
ProtoPNet + W $\times$ Act 45.70 $\pm$ 4.9 11.22$\pm$5.5 9.3-80.12 60.92 69.6 84.5
PiPNet(Covx) + Act 52.78$\pm$7.4 9.07$\pm$9.4 0.02-60.62 70.59 65.2 88.5
PiPNet(Covx) + W $\times$ Act 52.78 $\pm$ 7.4 9.07$\pm$9.4 10.83-95.66 70.59 65.2 88.5
PiPNet(Res50) + Act 42.40$\pm$17.5 17.4$\pm$2.4 0.18 - 1.21 63.19 60.8 86.46
PiPNet(Res50) + W $\times$ Act 42.40$\pm$17.5 17.4$\pm$2.4 9.43 - 93.15 63.19 60.8 86.46
Ours + Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 37.94 - 63.3 57.45 71.5 88.7
Ours + W $\times$ Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 59.33 - 91.73 57.45 71.5 88.7
Birds (CUB 200) Resnet50+Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 0.35 - 1.12 24.14 60.3 81.2
Resnet50+Rel $\times$ Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 2.43 - 5.73 24.14 60.3 81.2
ProtoPNet + Act 51.47$\pm$5.8 10.16$\pm$9.2 1.15-51.96 79.51 63.6 81.45
ProtoPNet + W $\times$ Act 51.47 $\pm$ 5.8 10.16$\pm$9.2 9.36-75.50 79.51 63.6 81.45
PiPNet(Covx) + Act 54.80$\pm$8.3 7.77$\pm$8.0 0.01-54.14 80.72 64.7 85.0
PiPNet(Covx) + W $\times$ Act 54.80 $\pm$ 8.3 7.77$\pm$8.0 5.13-92.83 80.72 64.7 85.0
PiPNet(Res50) + Act 42.28$\pm$3.2 15.57$\pm$15.7 0.14 - 2.35 77.83 65.9 82.0
PiPNet(Res50) + W $\times$ Act 42.28$\pm$3.2 15.57$\pm$15.7 9.40 - 89.77 77.83 65.9 82.0
PCMNet (Ours) + Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 41.82 - 75.24 57.45 67.1 84.7
PCMNet (Ours) + W $\times$ Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 55.13 - 89.88 57.45 67.8 84.7
@@
Table 1. Performance of PCMNet according to xAI metrics on the Stanford Cars and Birds datasets.
PCMNet is validated on standard benchmarks in prototype literature: Stanford Cars (Krause et al., 2013) (196 classes of different make and models) and CUB-200-2011 (Wah et al., 2011) (200 bird species, including attribute annotation for visual features). The ResNet50 model pretrained on ImageNet (Deng et al., 2009) is used as the feature backbone. Each training batch (32 images) is resized to 488×488, randomly cropped, erased, and padded. Optimization follows ADAM (Kingma and Ba, 2014) with a linear warm-up. We set regularization parameters $\lambda$ and $\gamma$ as same in (Oikarinen et al., 2023). $\alpha=1.5$, $\beta=2$, with $m_{1}=0.3$ and $m_{2}=1.5$ chosen via hyperparameter analysis in the suppl. material. More details on the experimental protocol are presented in the supplementary materials.
4.1. Explainability Metrics
To evaluate the explainability of final features, we rely on performance metrics established in the literature (Dreyer et al., 2024) with some minor modifications to express better comprehension.
@@ 4.1. Explainability Metrics
(a) Faithfulness:: Faithfulness(Dasgupta et al., 2022; Chan et al., 2022) evaluates how much the activated concepts influence the model’s final prediction, serving as a key metric for assessing the quality of explanations. One widely used approach for measuring faithfulness is concept deletion, which involves removing selected concepts from the model’s reasoning process and observing the resulting change in output confidence. Prior works, such as (Dreyer et al., 2024), typically assess this by removing only the single most important concept. However, this limited scope fails to capture the broader contribution of other activated concepts and may not reveal the full extent of their impact. To provide a more comprehensive assessment, we extend this approach by successively removing the top-$k$ most activated concepts and measuring the resulting drop in model confidence. A larger drop indicates higher faithfulness, as it suggests that the removed concepts were indeed critical to the model’s decision. This extended evaluation allows us to better quantify how much the explanation aligns with the model’s internal reasoning.
@@
(b) Stability: The extracted concepts from input images that share some semantic concepts must be stable and explainable for unseen images. To evaluate stability, we compute CAVs on $k$-fold subsets of the data ($k$ = 10 as default) similar to (Dreyer et al., 2024) for all classes at steps 2 and 3. We then map CAVs together using a Hungarian loss function and measure the cosine similarity between them.
@@
4.2. Comparison to Prototype xAI Methods
@@
To evaluate PCMNet, we report the classification accuracy and the XAI evaluation criteria discussed in Sec 4 and compare them to other ante-hoc methods ( ProtoPNet(Chen et al., 2019) and PIPNet(Nauta et al., 2023)) in tables 1. The consistency metric is reported in two ways: Intra and Inter, which are cosine similarities between activated concepts of the same class and different classes, respectively. Faithfulness is reported as F(n), measuring accuracy drops when the top n important concepts are removed from the decision. To compute the contribution of each concept or feature to the final decision, we evaluate two strategies: (1) using raw values of each concept (”Act”), and (2) computing a weighted contribution ( gradient) as the product of raw values and its corresponding weight in the classification layer (“W $\times$ Act”).
4.3. Ablation Studies
4.3.1. Number of Prototypes
We assess the impact of the number of parts on both classification accuracy and faithfulness in Table 2. As the model extracts more prototypical concepts, its decisions become more aligned with the underlying features, enhancing interpretability. Interestingly, when the number of prototypes is smaller, the faithfulness scores for the top concepts tend to be higher. This is because the model relies on a more compact set of prototypes, concentrating its decisions on fewer, more influential concepts.
K value Cars Birds
C Acc F(1)-F(5) C Acc F(1)-F(5)
0 (Resnet50) 84.1 1.54 - 6.61 83.2 1.21-5.56
4 83.5 79.96 - 91.01 83.5 71.42-85.64
5 84.1 70.11 – 91.33 84.9 66.7 - 85.9
6 85.2 66.02 – 90.65 85.7 59.9 - 86.7
7 86.7 62.47 – 91.25 86.0 55.7 - 87.3
8 87.5 59.33 – 91.73 85.1 53.0 - 86.4
9 86.4 55.10 – 91.52 84.4 52.5-86.1
Table 2. Impact on performance of the number of prototypes.
Figure 5. Visualizing activated concepts in the test set and examples from the training set for Stanford Cars and CUB datasets.
4.3.2. Effect of Modules
To evaluate the effectiveness of each component in PCMNet, we conduct an ablation study on the baseline PDiscoNet model, which extracts local parts without explicit concept learning. As shown in Table 3, we analyze classification accuracy (C Acc) and feature faithfulness under occlusion (F(3)) on both the Cars and Birds datasets, alongside the model size and FLOPs. Introducing the Marginal Clustering Center (MCC) loss into the baseline improves classification performance (from 84.2% to 86.1% on Cars) and more than doubles faithfulness scores, indicating that enforcing clustering consistency across part features supports more explainability. The inclusion of the Concept Mining Module (CMM)—responsible for extracting sparse and interpretable concepts—dramatically boosts faithfulness (from 4.9% to 55.6% on Birds), even with minimal impact on accuracy. When both MCC and CMM are combined, PCMNet achieves the best results across all metrics, including 87.5% classification accuracy and 67.4% faithfulness on Cars, with only a marginal increase in parameter count and computational cost. These findings underscore the complementary roles of MCC in enforcing cluster-level consistency and CMM in mining high-quality, human-aligned concepts.
@@ 4.3.2. Effect of Modules
C Acc F(3) C Acc F(3)
Baseline(van der Klis et al., 2023) 81.2 5.4 82.3 4.9 27.1M 17.2G
Base + MCC 86.1 9.6 84.5 9.1 27.2M 17.2G
Base + CMM 84.3 59.2 82.1 55.6 27.4M 17.5G
Base + MCC + CMM 87.5 67.4 84.7 64.9 27.5M 17.5G
PipPNet(Resnet) 86.4 62.7 82.0 58.6 25.1M 27.6G
PipPNet(Covx) 88.5 60.6 85.0 57.5 29.3M 32.3G
@@ C Acc F(3) C Acc F(3)
Table 3. Classification accuracy and faithfulness using different PCMNet modules.
Figure 6. t-SNE visualization of concept-level part features extracted by PCMNet. Colors represent different prototypes, while shapes indicate object classes.
4.4. Robustness to Occlusion
@@
To evaluate the robustness of PCMNet’s decision-making under occlusion, we conducted controlled experiments in which regions corresponding to the most activated concepts were systematically masked, and the resulting images were re-evaluated by the model. This experiment aims to assess the stability of model predictions when key interpretable regions are removed. Specifically, we identify the center of the most activated concept—based on the model’s response to clean images—using the concept representation (patch for PIPNet and ProtoPNet, mask for PCMNet), and then occlude a rectangular region centered around this point. We consider three levels of occlusion, masking 10%, 20%, and 30% of the image area. PCMNet’s performance under these conditions is compared against two other interpretable baselines: ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023).
Figure 7. Classification Accuracy and Faithfulness under Different Occlusion Levels.
Fig. 7 shows that PCMNet maintains stable accuracy at lower occlusion levels and degrades more gradually than ProtoPNet and PIPNet. This resilience stems from PCMNet’s broader part-based concepts, which introduce redundancy and enable accurate predictions even when key components are missing. Unlike patch-based models that rely on fixed, small regions, it extracts prototypes from semantically meaningful, wider regions. As a result, such methods experience sharp performance drops under occlusion due to their dependence on localized patterns, whereas PCMNet remains more robust by leveraging multiple related part concepts, preventing performance from relying too heavily on a single visual cue.
4.5. Qualitative Results
4.5.1. Concept Activation and Occlusion Robustness
@@
To further illustrate the interpretability and robustness of PCMNet, we compare the activated concepts of our method with PIPNet (Nauta et al., 2023) under clean and occluded conditions, as visualized in Fig. 5. For each method, we select two activated concepts on test images and retrieve training set examples with the highest activation values for those concepts. This comparison shows several key advantages of PCMNet:
@@
Broader Concept Coverage: PCMNet extracts more diverse and semantically rich concepts compared to PIPNet. Our method leverages attention-based masks instead of fixed patches to capture part-level semantics over a broader spatial context, leading to more varied and interpretable regions across different object parts.
More Informative Examples: The training examples retrieved by PCMNet for each concept exhibit more clearly activated and semantically aligned regions. This suggests that the discovered concepts better reflect the underlying factors influencing the model’s decision, thereby enhancing explanation quality.
@@
Robustness to Occlusion: Under occlusion, PIPNet often fails to highlight meaningful or consistent regions due to its reliance on limited regions. In contrast, PCMNet maintains interpretability by activating alternative, unoccluded parts that belong to the same or related concepts. This demonstrates the model’s ability to generalize concept reasoning even when certain visual cues are missing.
4.5.2. Visualization of Concept Prototypes
@@
To better understand the behavior of the learned prototypes in PCMNet, we visualize the extracted part-level features using t-SNE (van der Maaten and Hinton, 2008), as shown in Figure 6. Each point corresponds to a feature from a localized part, colors indicating the prototype (i.e., concept) and marker shapes denoting the object class. This visualization represents several insights:
@@
Disentanglement of Prototypes: Features assigned to different prototypes form well-separated clusters, suggesting that PCMNet effectively disentangles the latent interpretable part-based concept space. This separation implies that the Marginal Clustering Center loss and Concept Mining Module (CMM) encourages diverse and semantically distinct part concepts, which supports both interpretability of model decisions and robustness on occlusion.
@@
5. Conclusion
We presented PCMNet, a novel framework designed to bridge the gap between accuracy and interpretability in deep learning models. By mining part-prototypes from dynamic image regions, PCMNet improves upon existing patch-based prototype models by ensuring semantic coherence across instances. Our results show that PCMNet enhances faithfulness, concept stability, and explainability, outperforming existing interpretable models such as ProtoPNet and PIPNet. Moreover, our occlusion-based robustness analysis confirmed that PCMNet remains stable under missing or distorted input regions, demonstrating its reliability in real-world scenarios. Unlike conventional post-hoc explainability methods, PCMNet allows for more human-aligned, concept-based reasoning, making deep learning decisions more transparent and interpretable.
Appendix A Additional Experiment and Results
A.1. Implementation Details
We used Resnet50 as the backbone, which extracts 2048 features. We project these features into 512 features before global pooling, i.e., $d_{f}=512$. Based on our experiment and set of hyper-parameters, for the CUBs dataset, the $d_{c}=2954$ and Stanford Cars $d_{c}=3092$ are chosen. As the DBSCAN is used for clustering inside of each class, for each part and class, the number of clusters is not fixed.
The $\mathcal{L}_{\text{bl}}$(van der Klis et al., 2023) loss also is
(8) $\mathcal{L}_{\text{bl}}=\mathcal{L}_{\text{cond}}+\mathcal{L}_{\text{equi}}+\mathcal{L}_{\text{pres}},$

### v2

@@ Introduction
$\mathbf{z}^{i}=\{\mathcal{S}(\mathcal{C}(l,p,j),\mathbf{f}_{p}^{i})\;\forall p\in\{1..K\}\;\text{and}\;j\in\{1..L\}\},$ (6)
where $\mathbf{z}^{i}\in\mathbb{R}^{d_{c}}$ ans $\mathcal{S}$ is cosine distance. We aim to explain each output class of our model using a small set of interpretable concepts. We, therefore, train a sparse classifier on top of $z^{i}$ to obtain the final classification scores $o^{i}=W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b$ and the predicted class $\hat{y}^{i}=\arg\max(o^{i})$. Here, $W_{1}\in\mathbb{R}^{d_{c}\times L}$, $W_{2}\in\mathbb{R}^{d_{f}\times L}$ and $b\in\mathbb{R}^{L}$ denote the classification weights and bias term, respectively. This sparse layer is trained with the following sparse classification loss (wong2021leveraging):
$\mathcal{L}_{\text{cl}}=\mathcal{L}_{ce}(W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b,y^{i})+\lambda\mathcal{R}(W_{1}),$ (7)
where $\mathcal{L}_{\text{ce}}$ is the cross-entropy loss, $y^{i}$ is the label, $\lambda$ is the sparsity regularization strength and $\mathcal{R}(W)=(1-\gamma)\frac{1}{2}\|W\|_{F}+\gamma\|W\|_{1}$, here $\|W\|_{F}$ is the Forbenius norm and $\|W\|_{1}$ denotes the element-wise matrix.
Overall Training
We jointly optimize the network in an end-to-end manner:
$\mathcal{L}=\mathcal{L}_{\text{part}}+\beta\mathcal{L}_{\text{cl}},$ (8)
where $\beta$ is hyperparameter. We set $\beta=0$ during initial training and $\beta=2$ after learning part-prototypical concepts.
Experiments and Results
@@
Cars-196
Baseline 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 1.54 - 6.61 22.74 65.4 84.1
[…]
CUB-200
Resnet50 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 2.43 - 5.73 24.14 60.3 81.2
[…]
@@
Table 1: Performance of PCMNet according to xAI metrics on the Stanford Cars and Birds datasets.
Implementation Details
PCMNet is evaluated on two standard prototype-learning benchmarks: Stanford Cars (scar) (196 car classes) and CUB-200-2011 (cubs) (200 bird species). We use a pretrained ResNet50 as the backbone. Each training batch (32 images) is resized to 488×488 with random cropping and padding. Optimization is performed using Adam (kingma2014adam). Regularization parameters $\lambda$ and $\gamma$ follow (oikarinen2023label); we set $\alpha=1.5$, $\beta=2$, with margins $m_{1}=0.3$ and $m_{2}=1.5$ based on hyperparameter tuning analysis [^2] Full paper: https://arxiv.org/pdf/2504.12197.
End-to-End Multi-Stage Training: PCMNet modules (part discovery, concept clustering, and concept activation) are trained end-to-end in a unified pipeline over 40 epochs:
- Step 1 (initial epochs): The part discovery module is optimized to localize semantically meaningful part features. This stage continues until convergence of $\mathcal{L}_{\text{part}}$.
- Step 2 (intermediate epochs): Concept clustering runs intermittently (every 5 epochs) to reduce computation while keeping centroids up to date with changing representations.
- Step 3 (final epochs): The Concept Activation Mining Module (CMM) is trained alongside the backbone, using the updated centroids from Stage 2 to align activated features with discriminative concept prototypes. At inference, PCMNet behaves as a standard forward-pass classifier.
Explainability Metrics
To evaluate explainability, we use established metrics from (pcx), with minor adjustments.
@@
(a) Faithfulness: Faithfulness (fait1; fait2) evaluates how much the activated concepts influence the model’s final prediction, serving as a key metric for assessing the quality of explanations. One widely used approach for measuring faithfulness is concept deletion, which involves removing selected concepts from the model’s reasoning process and observing the resulting change in output confidence. Prior works, such as (pcx), typically assess this by removing only the single most important concept. However, this limited scope fails to capture the broader contribution of other activated concepts and may not reveal the full extent of their impact. To provide a more comprehensive assessment, we extend this approach by successively removing the top-$k$ most activated concepts and measuring the resulting drop in model confidence. A larger drop indicates higher faithfulness, as it suggests that the removed concepts were indeed critical to the model’s decision. This extended evaluation allows us to better quantify how much the explanation aligns with the model’s internal reasoning.
@@
(b) Stability: The extracted concepts from input images that share some semantic concepts must be stable and explainable for unseen images. To evaluate stability, we compute CAVs on $k$-fold subsets of the data ($k$ = 10 as default) similar to (pcx) for all classes at steps 2 and 3. We then map CAVs together using a Hungarian loss function and measure the cosine similarity between them.
@@
Comparison to Prototype xAI Methods
@@
To evaluate PCMNet, we report classification accuracy and XAI metrics, and compare them with other ante-hoc methods (ProtoPNet(ProtoPnet), PIPNet(PIPNet), ProtoViT (protoViT) and MCPNet (Wang2024MCPNet)) in Table 1. The consistency metric is reported in two forms: Intra and Inter, measuring the cosine similarity between activated concepts within the same class and across different classes, respectively. Faithfulness is reported as F(n), measuring accuracy drops when the top n important concepts are removed from the decision. To compute the contribution of each concept to the final decision, we compute a weighted score as the product of its raw value and its corresponding classification weight.
Ablation Studies
(a) Effect of Number of Concept Center centroids ($d_{c}$):
The size of the concept list ($d_{c}$) determines the semantic resolution of PCMNet and is influenced by both the number of parts and the number of training classes. As the number of parts ($K$) increases, so does the total number of extracted centroids. We analyze this dependency further in Appendix B2. While richer concept sets can improve interpretability, they also increase memory and computational cost.
To manage this, we apply a post-clustering refinement step that merges similar centroids within or across classes using hierarchical clustering (murtagh2011ward). Table 2 reports the effect of varying the merging level and threshold—defined as a percentage of the maximum inter-centroid distance—on the Cars dataset. We find that moderate merging (e.g., Level 1 with 10% threshold) significantly reduces $d_{c}$ with minimal impact on classification accuracy or faithfulness (F(3)). However, excessive merging (e.g., Level 3 at 10%) leads to performance degradation, suggesting the importance of controlled prototype compression.
Threshold (%) Level $d_{c}$ C Acc (%) F(3) (%)
0% 1 3092 90.2 67.4
2 2473 85.7 65.7
3 698 84.4 66.5
5% 1 2862 89.5 66.5
2 1805 83.6 66.2
3 645 80.5 65.7
10% 1 1657 86.2 67.9
2 1263 82.7 66.9
3 454 74.7 65.1
Table 2: Impact of hierarchical clustering thresholds and levels on the number of concept prototypes ($d_{c}$), classification accuracy, and interpretability (faithfulness F(3)) on the Cars dataset. Thresholds, expressed as a percentage of the maximum pairwise centroid distance.
(b) Effect of Modules:
To assess the impact of each PCMNet component, we conduct an ablation study on the baseline PDiscoNet model without concept learning. Additionally, the effect of using different types of concept vectors—prototypical and non-prototypical—is examined in Appendix B2. Table 3 reports classification accuracy (Acc) and faithfulness (F(3)) on the Cars and Birds datasets. Adding the Marginal Clustering Center (MCC) loss improves accuracy (e.g., from 81.2% to 87.8% on Cars) and doubles faithfulness, highlighting the benefits of enforcing cluster-level consistency. Incorporating the Concept Mining Module (CMM) yields a major boost in faithfulness (from 4.9% to 55.3% on Birds), with minimal cost to accuracy. Combining both modules yields the best overall performance: 90.2% accuracy and 67.4% faithfulness on Cars, with only +0.4M parameters and +0.3G FLOPs. Compared to recent interpretable baselines like PipPNet, ProtoPNet, ProtoViT, and MCPNet, PCMNet achieves a better balance between accuracy, faithfulness, and efficiency, especially outperforming ProtoViT in F(3) (+22.8%) and requiring fewer FLOPs (17.4G vs. 24.8G). This confirms the complementary value of MCC and CMM in delivering interpretable and efficient visual recognition. The trade-off information is discussed in the ArXiv version.
@@
Acc F(3) Acc F(3)
Baseline 81.2 5.4 82.3 4.9 27.1M 17.1G
Baseline + MCC 87.8 9.2 84.9 8.9 27.2M 17.1G
Baseline + CMM 85.8 58.6 82.5 55.3 27.4M 17.4G
PCMNet (Full) 90.2 67.4 85.1 64.9 27.5M 17.4G
PiPNet 86.4 62.7 82.0 58.6 25.1M 27.6G
@@
Table 3: Classification accuracy and faithfulness using different PCMNet modules.
Robustness to Occlusion
@@
To evaluate PCMNet robustness under occlusion, we conducted controlled experiments by masking regions linked to the most activated concepts and re-evaluating the modified images with the model. This experiment aims to assess the stability of model predictions when key interpretable regions are removed. Specifically, we identify the center of the most activated concept—based on the model’s response to clean images—using the concept representation (patch for PIPNet and ProtoPNet, mask for PCMNet), and then occlude a rectangular region centered around this point. We consider three levels of occlusion, masking 10%, 20%, and 30% of the image area. PCMNet’s performance under these conditions is compared against two other interpretable baselines: ProtoPNet (ProtoPnet) and PIPNet (PIPNet).
(a) Accuracy.
(b) Faithfulness F(3)
Figure 3: Classification accuracy and faithfulness for a growing level of occlusion.
Fig. 3.a shows that PCMNet degrades more gradually under occlusion compared to ProtoPNet and PIPNet. This robustness comes from PCMNet’s semantically rich and diverse part concepts, which capture complementary aspects of the object. Also, activated concepts from PCMNet preserve the faithfulness since they are extracted from not occluded regions as shown in Fig. 3.b and buttom of Fig. 4. By describing class-discriminative features from multiple perspectives, PCMNet enables more reliable predictions. In contrast, patch-based models depend heavily on fixed, localized cues and thus exhibit sharper performance drops.
Qualitative Results
Figure 4: visualizations of activated concept on test samples with matching training samples for Stanford Cars and CUB.
@@
To further illustrate the interpretability and robustness of PCMNet, we compare the activated concepts of our method with PIPNet under clean and occluded conditions, as visualized in Fig. 4. For each method, we select two activated concepts on test images and retrieve training set examples with the highest activation values for those concepts. This comparison shows several key advantages of PCMNet:
@@
Broader Concept Coverage: PCMNet extracts more diverse and semantically rich concepts compared to PIPNet. It leverages attention-based masks instead of fixed patches to capture part-level semantics over a broader spatial context, leading to more varied and interpretable regions across different object parts and improving explanation quality.
@@
Robustness to Occlusion: As shown at the bottom of Fig. 4, under occlusion, PIPNet often fails to highlight meaningful regions due to its reliance on limited regions. In contrast, PCMNet maintains interpretability by activating alternative, unoccluded parts that belong to the same or related concepts. This demonstrates the model’s ability to generalize concept reasoning even when certain visual cues are missing.
Figure 5: t-SNE visualization of concept-level part features extracted by PCMNet. Colors represent different prototypes, while shapes indicate object classes.
@@
To better understand the behavior of the learned prototypes in PCMNet, we visualize the extracted part-level features using t-SNE (tsne), as shown in Fig.  5. Each point corresponds to a feature from a localized part, colors indicating the prototype (i.e. concept) and marker shapes denoting the object class. This visualization represents several insights:
@@
Disentanglement of Prototypes: Features assigned to different prototypes form well-separated clusters, indicating that PCMNet effectively disentangles part-based concepts. This shows that the Marginal Clustering Center loss and CMM promote diverse, semantically distinct parts, enhancing interpretability and occlusion robustness.
@@
Conclusion
We presented PCMNet, a novel framework designed to bridge the gap between accuracy and interpretability in deep learning models. By mining part-prototypes from dynamic image regions, PCMNet improves upon existing patch-based prototype models by ensuring semantic coherence across instances. Our results show that PCMNet enhances faithfulness, concept stability, and explainability, outperforming existing interpretable models. Moreover, our occlusion-based robustness analysis confirmed that PCMNet remains stable under missing or distorted input regions, demonstrating its reliability in real-world scenarios. Unlike conventional post-hoc explainability methods, PCMNet allows for more human-aligned, concept-based reasoning, making deep learning decisions more transparent and interpretable.
Acknowledgment:
This research was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Digital Research Alliance of Canada.
Appendix A Supplementary: Additional Experiment and Results
Appendix A: Additional Implementation Details
In our model, $F^{i}=\mathcal{F}(x^{i})$ denotes the spatial feature map extracted by the backbone network $\mathcal{F}$, with shape $\mathbb{R}^{D\times w\times h}$, where $D$ is the backbone feature dimension, and $w\times h$ is the spatial resolution. The part-prototype module $\mathcal{P}$ produces a soft assignment mask $M^{i}=\mathcal{P}(F^{i})\in\mathbb{R}^{(K+1)\times w\times h}$, where each pixel location is softly assigned to one of the $K$ part and one background channels. Specifically, for a given 2D location $a=(a^{x},a^{y})$, the assignment scores across all parts are:
$M^{i}(a)=M^{i}[:,a^{x},a^{y}]\in\mathbb{R}^{K+1}.$ (9)
while the full mask for part $p$ is given by,
$M^{i}_{p}=M^{i}[p,:,:]\in\mathbb{R}^{w\times h}.$ (10)
We used Resnet50 as the backbone, which extracts 2048 features. We project these features into 512 features before global pooling, i.e., $d_{f}=512$. Based on our experiment and set of hyper-parameters, for the CUBs dataset, the $d_{c}=2954$ and Stanford Cars $d_{c}=3092$ are computed in stage 2. As the DBSCAN is used for clustering inside of each class, for each part and class, the number of clusters is not fixed.
Learning Curvature
Fig. 6 illustrates the training loss and validation accuracy of PCMNet over 40 epochs. In Step 1, only the backbone and part-discovery modules are trained using the part discrimination loss ($\mathcal{L}\text{part}$). Once this loss converges, Step 2 and 3 begin, during which concept centroids are extracted and the clustering loss ($\mathcal{L}\text{cl}$) is introduced, respectively. This causes a temporary spike in the loss, followed by a gradual decrease as the model adapts to reasoning over learned concepts. To keep the prototypes dynamic and data-aligned, centroid extraction and Step 2 are repeated every 5 epochs.
Figure 6: Training loss over 40 epochs showing three-stage learning dynamics.
Additional details of Baseline loss
The $\mathcal{L}_{\text{bl}}$(van2023pdisconet) loss also is
$\mathcal{L}_{\text{bl}}=\mathcal{L}_{\text{cond}}+\mathcal{L}_{\text{equi}}+\mathcal{L}_{\text{pres}},$ (11)